# A Communication Consistent Approach to Signal Temporal Logic Task Decomposition in Multi-Agent Systems

### v2

@@ 6.2 A decentralized solution
In (55), the superscript $t$ indicates the current iteration and is set on the upper-left to avoid confusion with the previous notation. The term $\rho_{rs}\in\mathbb{R}$ is a penalty variable constrained to be non-negative as per constraint (55b). The coefficient $c_{\rho}>0$ penalizes non-zero values of $\rho_{rs}$. At the same time, by (55c), each node $(r,s)$ holds the value at iteration $t$ of the additional consensus vectors $^{t}\bm{\lambda}_{rs,r^{\prime}s^{\prime}},\in\mathbb{R}^{(\sum_{(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}}m_{ij})},\;\forall(r^{\prime},s^{\prime})\in\mathcal{N}_{\Theta}((r,s))$ and receives the consensus vectors $^{t}\bm{\lambda}_{r^{\prime}s^{\prime},r,s}$ from the each neighbour $(r^{\prime},s^{\prime})\in\mathcal{N}_{\Theta}((r,s))$. The introduction of a consensus term like the one introduced in (55) is common in distributed optimization of a global optimization program with a separable structure as the one in (52) [29, 30]. Note that summing the left-hand side of constraint (55c) for each $(r,s)\in\mathcal{E}_{\pi}=\mathcal{V}_{\Theta}$ yields the original left-hand side in (52c). At this point Algorithm 1 borrowed from [29] applies to converge to an optimal solution of the vectors $\bm{\chi}_{rs},\;\forall(r,s)\in\mathcal{E}_{\pi}$, after sufficiently many iterations. While referring the reader to [29] for a detailed analysis of Alg. 1, we here seek to provide an intuitive explanation of the algorithm. Namely, at each iteration $t$, the edge $(r,s)\in\mathcal{V}_{\Theta}$ solves an instance of (55), where the penalty variable $\rho_{rs}\in\mathbb{R}$ relaxes the satisfaction of the shared constraint (55c). If we let $^{t}\rho_{rs}$ to be the optimal value of $\rho_{rs}$ at iteration $t$ and if we let $^{t}\bm{\mu}_{rs}\in\mathbb{R}^{\sum_{(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}}m_{ij}}$ to be the Lagrangian multiplier vector associated with (55c) at each iteration $t$, then the consensus vectors $\bm{\lambda}_{rs,r^{\prime}s^{\prime}}$ are updated by node $(r,s)$, at iteration $t$, as per line 7 in Alg, 1. Note that this update only requires the local knowledge of the Lagrangian multiplier vectors $\bm{\mu}_{rs}$, $\bm{\mu}_{r^{\prime}s^{\prime}},\;\forall(r^{\prime},s^{\prime})\in\mathcal{N}_{\Theta}((r,s))$. The update rate $^{t}\gamma>0$ must be known to all agents and such that $\lim_{t\rightarrow\infty}{}^{t}\gamma=\infty$ and $\lim_{t\rightarrow\infty}({}^{t}\gamma)^{2}<\infty$ to guarantee convergence [29]. If we let $\bm{\chi}_{rs}^{\star}]_{(r,s)\in\mathcal{E}_{\pi}}$ to be a non-unique optimal (feasible) solution of (52), then, by [29, Thm II.6], it is known that the iterates of optimal solutions $\{[{}^{t}\bm{\chi}_{rs}]_{\forall(r,s)\in\mathcal{E}_{\pi}}\}$ found by Alg. 1 converge asymptotically to one such solution $\bm{\chi}_{rs}^{\star}]_{(r,s)\in\mathcal{E}_{\pi}}$. This, in turns, entails that $\lim_{t\rightarrow\infty}{}^{t}\rho_{rs}=0,\;\forall(r,s)\in\mathcal{V}_{\Theta}$. Once Alg. 1 terminates for all the agents, a set of optimal parameters $(\bm{\eta}_{rs}^{l})^{\star},\;\forall l\in\bar{\mathcal{K}}_{rs}$, is obtained from each edge $(r,s)\in\mathcal{E}_{\pi}$ such that the parametric tasks $\bar{\varphi}_{rs}^{ij}\equiv\varphi_{rs}^{l},\;\forall l\in\bar{\mathcal{K}}_{rs}$ are fully determined as well as the new global task $\bar{\psi}$ as per (20), thus solving Problem 1.
@@
2:initialize $^{0}\bm{\lambda}_{rs,r^{\prime}s^{\prime}},\;\forall(r^{\prime}s^{\prime})\in\mathcal{N}_{\Theta}((r,s))$ and $c_{\rho}>>0$
@@
5:  get $\left(\left({}^{t+1}\bm{\chi}_{rs},{}^{t+1}\rho_{rs}\right),{}^{t+1}\bm{\mu}_{rs}\right)$ solution of (55)
@@
8:  $t=t+1$
With these considerations, our final main result is presented next.
Theorem 6.15.
@@
When (50) is feasible, then the satisfaction of the constraints (50b)-(50c) implies, by Prop. 5.4, 5.6, 5.8 and 5.10, that the resulting tasks $\phi_{ij},\,\forall(i,j)\in\mathcal{E}_{\bar{\psi}}$ are communication consistent and are not conflicting conjunctions. Moreover, by the conditions of Lemma 1, which are enforced via constraint (50d), we know that for each inconsistent task $\phi_{ij}$ with $(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$ in $\mathcal{G}_{\psi}$, the conjunction of tasks $\bar{\phi}^{ij}=\land_{(r,s)\in\epsilon(\bm{\pi}_{i}^{j})}\bar{\varphi}_{rs}^{ij}$ as per (21) is such that $(\bm{x}(t),0)\models\bar{\phi}^{ij}\Rightarrow(\bm{x}(t),0)\models\phi_{ij}$. From these considerations, the result follows.
6.3 Computational aspects
The optimization problem (50) has a compact convex domain and a convex objective such that a bounded set of solutions to (50) exists if the domain set is not empty. The infeasibility of (50) occurs when the intersection between the feasible set of (50c) and (50d) is empty. This intuitively happens when an edge $(r,s)\in\mathcal{E}_{\pi}$ is overloaded with an excessive amount of tasks including conflicting conjunctions that can not be resolved. When (50) is distributed among the agents as per (52), the convergence to the solution of the centralized problem is sublinear [29] and the dimensionality of (52) (in terms of number of variables) scales as $(n+1)|\mathfrak{Q}_{rs}|+n|\mathcal{K}_{rs}|$, where $|\mathcal{K}_{rs}|$ is the number of tasks decomposed over the edge $(r,s)$ and $|\mathfrak{Q}_{rs}|$ is the number of conflicting conjunctions over the edge. Intuitively, the complexity of (52) grows when edges are overloaded with many tasks that are potentially in conflict. On the other hand, if we let $\bar{m}$ be the maximum number of rows among the matrices $M_{ij}$ in (50d), then the number of exchanged Lagrangian multiplier variables ( see Alg. 1, line 4) has a worst case scaling of $\bar{m}\cdot|\mathcal{K}_{rs}|$. Thus, the more complex the polytope representation of the predicate level sets $\mathcal{B}_{ij}$ is, the more information has to be exchanged.
@@ 6.3 Computational aspects
We highlight that while the edge-computing graph is a convenient abstraction for our decomposition, either agent $r\in\mathcal{V}$ or $s\in\mathcal{V}$ will be effectively solving (55) over the edge $(r,s)\in\mathcal{E}_{\pi}=\mathcal{V}_{\Theta}$. Since the communication graph $\mathcal{G}_{c}$ is acyclic and $\mathcal{E}_{\pi}\subseteq\mathcal{E}_{c}$, then it is possible to select the computing nodes by selecting the leaf nodes of $\mathcal{G}_{c}$ as computing and iteratively select computing nodes along the paths from the leaf agent to the root of the graph. In this way the case in which a single agent $r$ solves an instance of (55) for both the edges $(r,s^{\prime})$ and $(r,s^{\prime\prime})$ for some $((r,s^{\prime}),(r,s^{\prime\prime}))\in\mathcal{E}_{\Theta}$, is avoided. Moreover, consider again two neighbouring edges $((s^{\prime},r),(r,s^{\prime\prime})\in\mathcal{E}_{\Theta}$ over the computing graph $\mathcal{G}_{\Theta}$ such that the node $r$ is shared. Then exchanging the consensus vectors and Lagrangian parameters $^{t}\bm{\mu}_{rs^{\prime}},\;{}^{t}\bm{\mu}_{rs^{\prime\prime}},\;{}^{t}\mathcal{\bm{\lambda}}_{rs^{\prime},rs^{\prime\prime}}$ and $^{t}\mathcal{\bm{\lambda}}_{rs^{\prime\prime},rs^{\prime}}$ among the neighbouring edges $(s^{\prime},r)$ and $(r,s^{\prime\prime})$ over the communication graph $\mathcal{G}_{c}$, takes either 1 communication hop (if $s^{\prime}$ and $r$ are the computing agents for $(s^{\prime},r)$ and $(r,s^{\prime\prime})$, respectively), or 2 communication hops (if $s^{\prime}$ and $s^{\prime\prime}$ are the computing agents for $(s^{\prime},r)$ and $(r,s^{\prime\prime})$, respectively).
7 Simulations
Table 1: Combined Edge-computing graph data
@@ 7 Simulations
Figure 6: (a) Optimal penalty values $^{t}\rho_{rs}\;\forall(r,s)\in\mathcal{E}_{\pi}$ and decomposition accuracy $\sum_{l\in\bar{\mathcal{K}}_{rs}}\alpha^{l}_{rs}$ as per Alg. 1. (b) Agent’s trajectories in the exploration phase (left) and return phase (right). Black arrows represent the direction of movement. A cyan square represents the main base. Red circles represent interest points to be visited.
@@
We apply the developed algorithm for a space exploration task. A team of 15 agents with single integrator dynamics $\dot{\bm{x}}_{i}=\bm{u}_{i}\in\mathbb{R}^{2},\forall i=1,\ldots 15$ is considered for the mission, where $\bm{x}_{i}$ represents the position of each agent and such that $\|\bm{u}_{i}\|\leq 1.8\;km/h$. The exploration consists of taking three panoramic images of the Thira crater on Mars’ southern hemisphere from three different sides of the crater and visiting smaller craters around Thira. The interest points are represented by red solid circles in Figure 6, with Thira being the largest and marked by a yellow dot. A cyan square represents the location of the main base station from which the agents initially depart from to accomplish the mission. Each agent is assumed to have a maximum communication range of $r_{c}=8.5$ km. The mission consists of a first exploration phase from the base station located at $[0,0]$ to the interest points, which lasts 20 hours. During the exploration phase, agent 1 can be considered the global leader of the team and the one holding the information about the interest points location such that the other agents 2-15 can reach their target location by staying in formation with agent 1. This phase is expressed by the taks $G_{[10,20]}\langle{}^{4}A_{1}(\bm{x}_{1}-[5\;0]^{T})-\bm{1}_{5}\cdot 0.4\rangle$, $F_{[10,15]}\langle{}^{6}A_{1,8}(\bm{e}_{1,8}-[10\;4]^{T})-\bm{1}_{6}\cdot 0.4\rangle$, $F_{[10,15]}\langle{}^{6}A_{1,4}(\bm{e}_{1,4}-[10\;-4]^{T})-\bm{1}_{6}\cdot 0.4\rangle$, $G_{[13,15]}\langle{}^{4}A_{1,14}(\bm{e}_{1,14}-[-12\;5.5]^{T})-\bm{1}_{4}\cdot 0.7\rangle$, $F_{[10,15]}\langle{}^{5}A_{1,13}(\bm{e}_{1,13}-[-6\;-6]^{T})-\bm{1}_{5}\cdot 0.4\rangle$, $F_{[13,15]}\langle{}^{5}A_{1,5}(\bm{e}_{1,5}-[-3.5\;-4]^{T})-\bm{1}_{5}\cdot 0.4\rangle$, $G_{[10,15]}\langle{}^{5}A_{1,6}(\bm{e}_{1,6}-[0\;2]^{T})-\bm{1}_{5}\cdot 0.4\rangle$, $G_{[10,15]}\langle{}^{5}A_{1,2}(\bm{e}_{1,2}-[0\;-2]^{T})-\bm{1}_{5}\cdot 0.4\rangle$, $G_{[10,15]}\langle{}^{5}A_{1,11}(\bm{e}_{1,11}-[-4\;0]^{T})-\bm{1}_{5}\cdot 0.4\rangle$, $G_{[18,20]}\langle{}^{5}A_{10,15}(\bm{e}_{10,15}-[0\;3]^{T})-\bm{1}_{5}\cdot 0.3\rangle$. The notation $^{n}A_{ij}$ indicates the matrix $A_{ij}$ associated with a regular polytope with $n$ sides i.e for $n=5$ then $^{n}A_{ij}$ is the matrix with 5 rows, where each row is the normal vector to the side of a regular pentagon. It is not necessary for our approach that the polytopes $\mathcal{P}$ have any regularity. Informally, the conjunction of the previous tasks implies that agents 1, 8 and 4 are involved in taking three panoramic images of Thira, with agent 1 (red trajectory in Fig. 6) approaching Thira from the west side leaving from the base station, while agents 4 and 8 reach the southern and northern sides of the crater respectively by maintaining an angle of 45 degrees with agents 1, thus reaching a triangular formation around Thira. Agent 14 visits the upper-left crater (marked as #A\#A), while agents 13 and 5 visit the lower craters (#B\#B ad #C\#C respectively). At the same time, agents 11, 6 and 2 should achieve a triangular formation around 1. Lastly, agent 15 and 10 should remain within 200 m from each other. After the exploration phase, the agents should group again around agent 1 in a radius of 2 km, while agent 1 returns to the base station effectively “dragging” the team back to base. This phase is described via the tasks $G_{[30,35]}\langle{}^{8}A_{i}(\bm{x}_{1}-[0\;0]^{T})-\bm{1}_{8}\cdot 0.2\rangle$, $G_{[30,40]}\langle{}^{8}A_{1,i}(\bm{e}_{1,i}-[0\;0]^{T})-\bm{1}_{8}\cdot 2\rangle,\forall i=2,\ldots 15$.
Figure 6 shows the exploration phase (left panel) from time 0h to 20h, and the return phase (right panel) from time 20h to 40h. The task graph ($\mathcal{G}_{\psi}$), communication graph ($\mathcal{G}_{c}$) and the new task graph ($\mathcal{G}_{\bar{\psi}}$) obtained implementing the distributed task decomposition algorithm presented in Section 6.2 are shown in Figure 7. To achieve the task satisfaction, the agents are controlled via the Control Barrier Functions-based controller proposed in [40], exploiting only the state information of the neighbours in $\mathcal{N}_{c}(i)$ while enforcing collision avoidance. The controller in [40] requires input-affine nonlinear dynamics, but more general dynamics can be considered, as long as a valid controller leveraging only information for the agents in $\mathcal{N}_{c}(i)$ is designed to satisfy $\bar{\psi}$. As of now, feedback controllers that can handle general nonlinear dynamics, STL constraints, collision avoidance, and communication constraints simultaneously are not widely studied and are an active area of research.
@@
The edge computing graph $\mathcal{G}_{\Theta}(\mathcal{V}_{\Theta},\mathcal{E}_{\Theta})$ is not reported due to space limitations, but it can be computed recalling that $\mathcal{V}_{\Theta}=\mathcal{E}_{\pi}$ and $\mathcal{E}_{\Theta}$ is as per (54). The task graph $\mathcal{G}_{\bar{\psi}}$ is computed before the start of the mission at time 0h, where Alg. 1 is run for 3500 iterations by each agent in $\mathcal{G}_{\Theta}$ with an average computational time of 52.133s/agent52.133\,s/agent (assuming no delay in exchanging the consensus and Lagrangian multiplier variables). In Table 1 we list the number of decomposition paths passing through a specific edge $(r,s)\in\mathcal{E}_{\pi}$ ($|\Pi_{rs}|$), the dimension of the variables $\bm{\chi}_{rs}$, the total number of shared constraints ($\sum_{(i,j)\in\Pi_{rs}}m_{ij}$) and the number of conflicting conjunctions sets of tasks ($|\mathfrak{Q}_{rs}|$). Figure 6(a) shows the evolution of the penalties values $^{t}\rho_{r,s}\;\forall(r,s)\in\mathcal{E}_{\pi}$ and the accuracy of decomposition for each task over a range of 3500 iterations. In particular, the time evolution of the optimal penalties $^{t}\rho_{rs}$ indicates that a valid task decomposition is found after 600 iterations of Alg. 1, while the decomposition accuracy as per (35) keeps improving before stabilizing at 3500 iterations. Since the optimization program (55) is only quasi-convex, the solution of the task decomposition is not unique. We used the open-source library CasADi to solve (55), leveraging an interior-point method solver on an Intel-Core i7-1265U .
@@
Figure 7: (Upper-left) original task graph $\mathcal{G}_{\psi}$, (upper-right) new task graph $\mathcal{G}_{\bar{\psi}}$, (lower-left) edge-computing graph $\mathcal{G}_{\Theta}$ and (lower-right) communication graph $\mathcal{G}_{c}$.
8 Conclusions

### v3

@@ VI-B A decentralized solution
In (55), the superscript $t$ indicates the current iteration and is set on the upper-left to avoid confusion with the previous notation. The term $\rho_{rs}\in\mathbb{R}$ is a penalty variable constrained to be non-negative as per constraint (55b). The coefficient $c_{\rho}>0$ penalizes non-zero values of $\rho_{rs}$. At the same time, by (55c), each node $(r,s)$ holds the value at iteration $t$ of the additional consensus vectors $^{t}\bm{\lambda}_{rs,r^{\prime}s^{\prime}},\in\mathbb{R}^{(\sum_{(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}}m_{ij})},\;\forall(r^{\prime},s^{\prime})\in\mathcal{N}_{\Theta}((r,s))$ and receives the consensus vectors $^{t}\bm{\lambda}_{r^{\prime}s^{\prime},r,s}$ from the each neighbour $(r^{\prime},s^{\prime})\in\mathcal{N}_{\Theta}((r,s))$. The introduction of a consensus term like the one introduced in (55) is common in distributed optimization of a global optimization program with a separable structure as the one in (53) [29, 30]. Note that summing the left-hand side of constraint (55c) for each $(r,s)\in\mathcal{E}_{\pi}=\mathcal{V}_{\Theta}$ yields the original left-hand side in (53c). At this point Algorithm 1 borrowed from [29] applies to converge to an optimal solution of the vectors $\bm{\chi}_{rs},\;\forall(r,s)\in\mathcal{E}_{\pi}$, after sufficiently many iterations. While referring the reader to [29] for a detailed analysis of Alg. 1, we here seek to provide an intuitive explanation of the algorithm. Namely, at each iteration $t$, the edge $(r,s)\in\mathcal{V}_{\Theta}$ solves an instance of (55), where the penalty variable $\rho_{rs}\in\mathbb{R}$ relaxes the satisfaction of the shared constraint (55c). If we let $^{t}\rho_{rs}$ to be the optimal value of $\rho_{rs}$ at iteration $t$ and if we let $^{t}\bm{\mu}_{rs}\in\mathbb{R}^{\sum_{(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}}m_{ij}}$ to be the Lagrangian multiplier vector associated with (55c) at each iteration $t$, then the consensus vectors $\mathcal{\lambda}_{rs,r^{\prime}s^{\prime}}$ are updated by node $(r,s)$, at iteration $t$, as per line 7 in Alg, 1. Note that this update only requires the local knowledge of the Lagrangian multiplier vectors $\bm{\mu}_{rs}$, $\bm{\mu}_{r^{\prime}s^{\prime}},\;\forall(r^{\prime},s^{\prime})\in\mathcal{N}_{\Theta}((r,s))$. The update rate $^{t}\gamma>0$ is must be known to all agents and such that $\lim_{t\rightarrow\infty}{}^{t}\gamma=\infty$ and $\lim_{t\rightarrow\infty}({}^{t}\gamma)^{2}<\infty$ to guarantee convergence [29]. If we let $\bm{\chi}_{rs}^{\star}]_{(r,s)\in\mathcal{E}_{\pi}}$ to be a non-unique optimal (feasible) solution of (53), then, by [29, Thm II.6], it is known that the iterates of optimal solutions $\{[{}^{t}\bm{\chi}_{rs}]_{\forall(r,s)\in\mathcal{E}_{\pi}}\}$ found by Alg. 1 converge asymptotically to one such solution $\bm{\chi}_{rs}^{\star}]_{(r,s)\in\mathcal{E}_{\pi}}$. This, in turns, entails that $\lim_{t\rightarrow\infty}{}^{t}\rho_{rs}=0,\;\forall(r,s)\in\mathcal{V}_{\Theta}$. Once Alg. 1 terminates for all the agents, a set of optimal parameters $(\bm{\eta}_{rs}^{l})^{\star}\;\forall l\in\bar{\mathcal{K}}_{rs}$, is obtained from each edge $(r,s)\in\mathcal{E}_{\pi}$ such that the parametric tasks $\varphi_{rs}^{l}\underset{y_{\Pi_{rs}}}{=}\bar{\varphi}_{rs}^{ij}$ as per (32) are fully determined as well as the new global task $\bar{\psi}$ as per (21), thus solving Problem 1.
@@
2:initialize $^{0}\bm{\lambda}_{rs,r^{\prime}s^{\prime}}$ arbitrarily $\forall(r^{\prime}s^{\prime})\in\mathcal{N}_{\Theta}((r,s))$ and $c_{\rho}>>0$
@@
5:  compute $\left(\left({}^{t+1}\bm{\chi}_{rs},{}^{t+1}\rho_{rs}\right),{}^{t+1}\bm{\mu}_{rs}\right)$ as primal-dual
                 optimal of (55)
@@
8:  end for
9:  $t=t+1$
10:end while
With these considerations, the final main result of our decomposition is presented next.
Theorem 1
@@
When (51) is feasible, then the satisfaction of the constraints (51b)-(51c) implies, by Prop. 6, 7, 8 and 9, that the resulting tasks $\phi_{ij},\,\forall(i,j)\in\mathcal{E}_{\bar{\psi}}$ are communication consistent and are not conflicting conjunctions. Moreover, by the conditions of Lemma 1, which are enforced via constraint (51d), we know that for each inconsistent task $\phi_{ij}$ with $(i,j)\in\mathcal{E}_{\psi}\setminus\mathcal{E}_{c}$, the conjunction of tasks $\bar{\phi}^{ij}=\land_{(r,s)\in\epsilon(\bm{\pi}_{i}^{j})}\bar{\varphi}_{rs}^{ij}$ as per (22) is such that $(\bm{x}(t),0)\models\bar{\phi}^{ij}\Rightarrow(\bm{x}(t),0)\models\phi_{ij}$. From these considerations, the result of the theorem follows. ∎
VI-C Computational aspects
Concerning some computational aspects, the optimization problem (51) has a compact convex domain and convex objective such that a bounded set of solutions to (51) exists if the feasible domain set is not empty. Hence, the proposed decomposition process is sound. The infeasibility of (51) occurs when the intersection between the domain of (51c) and (51d) is empty. This can be intuitively understood as one parametric polytope $\mathcal{B}_{rs}^{l}=\mathcal{P}(A_{rs}^{l},\bm{c}_{rs}^{l},\alpha_{rs}^{l}\bm{z}^{l}_{rs})$, for some $l\in\bar{\mathcal{K}}_{rs}$, being constrained with an excessive number of intersections that induce a scale factor $\alpha_{rs}^{l}>1$, thus impeding the satisfaction of the shared constraint (51d) due to the result in Prop. 5. To resolve these pathological cases, it is possible to change the structure of the communication topology by leveraging the communication tokens $q_{ij}^{c}$ hoping that a different selection of decomposition edges $\mathcal{E}_{\pi}$ provides a feasible solution, but we do not provide a direct solution in this direction, which we leave as future work. A few final comments regarding the way information is exchanged over the edge-computing graph $\mathcal{G}_{\Theta}$ and the amount of information required to run Algorithm 1 are given.
@@ VI-C Computational aspects
First note that that while the edge-computing graph is a convenient abstraction for our decomposition, either agent $r\in\mathcal{V}$ or $s\in\mathcal{V}$ will be effectively solving (55). Since the communication graph $\mathcal{G}_{c}$ is acyclic and $\mathcal{E}_{\pi}\subseteq\mathcal{E}_{c}$, then it is possible to select the computing nodes by selecting the leaf nodes of $\mathcal{G}_{c}$ as computing and iteratively select computing nodes along the paths from the leaf agent to the root of the graph. In this way the case in which a single agent $r$ solves an instance of (55) for both the edges $(r,s^{\prime})$ and $(r,s^{\prime\prime})$ for some $((r,s^{\prime}),(r,s^{\prime\prime}))\in\mathcal{E}_{\Theta}$ is avoided. Moreover, consider again two neighbouring edges $((s^{\prime},r),(r,s^{\prime\prime})\in\mathcal{E}_{\Theta}$ over the computing graph $\mathcal{G}_{\Theta}$ such that the node $r$ is shared. Then exchanging the consensus vectors and Lagrangian parameters $^{t}\bm{\mu}_{rs^{\prime}},\;{}^{t}\bm{\mu}_{rs^{\prime\prime}},\;{}^{t}\mathcal{\bm{\lambda}}_{rs^{\prime},rs^{\prime\prime}}$ and $^{t}\mathcal{\bm{\lambda}}_{rs^{\prime\prime},rs^{\prime}}$ among the neighbouring edges $(s^{\prime},r)$ and $(r,s^{\prime\prime})$ over the communication graph $\mathcal{G}_{c}$, takes either 1 communication hop (if $s^{\prime}$ and $r$ are the computing agents for $(s^{\prime},r)$ and $(r,s^{\prime\prime})$, respectively), or 2 communication hops (if $s^{\prime}$ and $s^{\prime\prime}$ are the computing agents for the edge $(s^{\prime},r)$ and $(r,s^{\prime\prime})$, respectively).
VII Simulations
Task Operator Predicate
Exploration task
$\phi_{1}$ $G_{[10,20]}$ $\scriptstyle\langle{}^{4}A_{1}\left(\bm{x}_{1}-[\scriptscriptstyle 5\;\scriptscriptstyle 0]^{T}\right)-\bm{1}_{5}\cdot 0.4\rangle$
$\phi_{1,8}$ $F_{[10,15]}$ $\scriptstyle\langle{}^{6}A_{1,8}\left(\bm{e}_{1,8}-[\scriptscriptstyle 10\;\scriptscriptstyle 4]^{T}\right)-\bm{1}_{6}\cdot 0.4\rangle$
$\phi_{1,4}$ $F_{[10,15]}$ $\scriptstyle\langle{}^{6}A_{1,4}\left(\bm{e}_{1,4}-[\scriptscriptstyle 10\;\scriptscriptstyle-4]^{T}\right)-\bm{1}_{6}\cdot 0.4\rangle$
$\phi_{1,14}$ $G_{[13,15]}$ $\scriptstyle\langle{}^{4}A_{1,14}\left(\bm{e}_{1,14}-[\scriptscriptstyle-12\;\scriptscriptstyle 5.5]^{T}\right)-\bm{1}_{4}\cdot 0.7\rangle$
$\phi_{1,13}$ $F_{[10,15]}$ $\scriptstyle\langle{}^{5}A_{1,13}\left(\bm{e}_{1,13}-[\scriptscriptstyle-6\;\scriptscriptstyle-6]^{T}\right)-\bm{1}_{5}\cdot 0.4\rangle$
$\phi_{1,5}$ $F_{[13,15]}$ $\scriptstyle\langle{}^{5}A_{1,5}\left(\bm{e}_{1,5}-[\scriptscriptstyle-3.5\;\scriptscriptstyle-4]^{T}\right)-\bm{1}_{5}\cdot 0.4\rangle$
$\phi_{1,6}$ $G_{[10,15]}$ $\scriptstyle\langle{}^{5}A_{1,6}\left(\bm{e}_{1,6}-[\scriptscriptstyle 0\;\scriptscriptstyle 2]^{T}\right)-\bm{1}_{5}\cdot 0.4\rangle$
$\phi_{1,2}$ $G_{[10,15]}$ $\scriptstyle\langle{}^{5}A_{1,2}\left(\bm{e}_{1,2}-[\scriptscriptstyle 0\;\scriptscriptstyle-2]^{T}\right)-\bm{1}_{5}\cdot 0.4\rangle$
$\phi_{1,11}$ $G_{[10,15]}$ $\scriptstyle\langle{}^{5}A_{1,11}\left(\bm{e}_{1,11}-[\scriptscriptstyle-4\;\scriptscriptstyle 0]^{T}\right)-\bm{1}_{5}\cdot 0.4\rangle$
$\phi_{10,15}$ $G_{[18,20]}$ $\scriptstyle\langle{}^{5}A_{10,15}\left(\bm{e}_{10,15}-[\scriptscriptstyle 0\;\scriptscriptstyle 3]^{T}\right)-\bm{1}_{5}\cdot 0.3\rangle$
Return task
$\phi_{1}$ $G_{[30,35]}$ $\scriptstyle\langle{}^{8}A_{i}\left(\bm{x}_{1}-[\scriptscriptstyle 0\;\scriptscriptstyle 0]^{T}\right)-\bm{1}_{8}\cdot 0.2\rangle$
$\phi_{1,i}$ $G_{[30,40]}$ $\scriptstyle\langle{}^{8}A_{1,i}\left(\bm{e}_{1,i}-[\scriptscriptstyle 0\;\scriptscriptstyle 0]^{T}\right)-\bm{1}_{8}\cdot 2\rangle,\forall i=2,\ldots 15$
TABLE I: List of tasks for the exploration and return phase.
TABLE II: Combined Edge-computing graph data
@@ VII Simulations
(a) Evolution of penalties optimal values $^{t}\rho_{rs}\;\forall(r,s)\in\mathcal{E}_{\pi}$ as per Alg. 1
(b) (Panel 0): Exploration phase over the interval 0-20h. (Panel 1): Snapshot of the achieved configuration at time 20h. (Panel 2) Return phase from time 20-40h. (Panel 3) Snapshot of the achieved configuration at time 40h.
Figure 6:
@@
We apply the developed algorithm for a space exploration task. A team of 15 agents with single integrator dynamics $\dot{\bm{x}}_{i}=\bm{u}_{i}\in\mathbb{R}^{2},\forall i=1,\ldots 15$ is considered for the mission, where $\bm{x}_{i}$ represents the position of each agent and such that $\|\bm{u}_{i}\|\leq 1.8\;km/h$. The exploration consists of taking three panoramic images of the Thira crater on Mars’ southern hemisphere from three different sides of the crater and visiting smaller craters around Thira. The interest points are represented by red dashed circles in Figure 6(b), with Thira being the largest and marked by a yellow dot. A cyan square represents the location of the main base station from which the agents initially depart to accomplish the mission. Each agent is assumed to have a maximum communication range of $r_{c}=8.5$ km. The mission consists of a first exploration phase from the base station located at $[0,0]$ to the interest points, which lasts 20 hours. During the exploration phase, agent 1 can be considered the global leader of the team and the one holding the information about the interest points location such that the other agents 2-15 can reach their target location by staying in formation with agent 1 according to the collaborative tasks in Table I. Specifically, agents 1,8 and 4 take charge of taking the three panoramic images of Thira, with agent 1 (red trajectory in Fig. 6(b)) approaching Thira from the west side leaving from the base station, while agents 4 and 8 reach the southern and northern sides of the crater respectively by maintaining an angle of 45 degrees with agents 1, thus reaching a triangular formation around Thira. Agent 14 visits the upper-left crater (marked as #A\#A), while agents 13 and 5 visit the lower craters (#B\#B ad #C\#C respectively). At the same time, agents 11,6 and 2 should achieve a triangular formation around 1. Lastly, agent 15 and 10 should remain within 200 m from each other. After the exploration phase, the agents should group again around agent 1 in a radius of 2 km, while agent 1 returns to the base station effectively “dragging” the team back to base. The list of tasks for the exploration and return phase are shown in Table I, where $\mathcal{P}({}^{n}A_{ij},\bm{c}_{ij},\bm{1}_{n}\beta_{ij})$ represents a regular polytope with $n$ sides, with each side having distance $\beta_{ij}>0$ from the centre $\bm{c}_{ij}$ (i.e $n=5$ represents a regular penthagon). It is not an assumption of our approach that the polytopes $\mathcal{P}$ should have any regularity. The agents are controlled via the Control Barrier Functions-based controller proposed in [42], such that each agent only exploits the state information of its communication neighbours in $\mathcal{N}_{c}$, where the communication graph is represented in Figure 7. Figure 6(b) shows the exploration phase on the two upper panels, where panel 0 represent the exploration from time 0h to 20h, while panel 1 represents a snapshot of the final configuration at time 20h. The return phase is represented in panels 2 and 3, where panel 2 shows the trajectories of the agents from time 20h to 40h, while panel 3 shows a snapshot of the final configuration at time 40h. Pink polytopes and black arrows apply to represent the original truth set of the tasks in Table I. The task graph ($\mathcal{G}_{\psi}$), communication graph ($\mathcal{G}_{c}$) and computing graph ($\mathcal{G}_{\Theta}$) are represented in Figure 7 together with the resulting task graph $\mathcal{G}_{\bar{\psi}}$ obtained implementing the distributed task decomposition algorithm presented in Section VI-B.
@@
The task decomposition is computed before the start of the mission at time 0h, where Alg 1 is run for 3500 iterations by each agent in $\mathcal{G}_{\theta}$ with an average computational time of $52.133s$ (assuming no delay in exchanging the consensus and lagrangian multiplier variables). General information about the number of decomposition paths passing through a specific edge $(r,s)\in\mathcal{E}_{\pi}$ ($|\Pi_{rs}|$), the dimension of the variables $\bm{\chi}_{rs}$, the total number of shared constraints ($\sum_{(i,j)\in\Pi_{rs}}m_{ij}$) and the number of conflicting conjunctions sets of tasks ($|\mathfrak{Q}_{rs}|$) are given in Table II. Figure 6(a) shows the evolution of the penalties values $^{t}\rho_{r,s}\;\forall(r,s)\in\mathcal{E}_{\pi}$ and the accuracy of decomposition for each task over a range of 3500 iterations. In particular, the time evolution of the optimal penalties $^{t}\rho_{rs}$ indicates that a valid task decomposition is found after 600 iterations of Alg. 1, while the decomposition accuracy as per (36) keeps improving before stabilizing at 3500 iterations. We note that since the optimization program (55) is only quasi-convex, the solution of the task decomposition is not unique. We used the open-source library CasADi ([43]) to solve the optimization program (55) leveraging an interior-point method solver on an Intel-Core i7-1265U. We highlight that agents in Figure 6 only exploit information from neighbouring agents in $\mathcal{G}_{c}$ for both controlling their state in Figure 6(b) and achieving the task decomposition before staring the mission.
@@
Figure 7: (Upper-left) original task graph $\mathcal{G}_{\psi}$, (upper-right) new task graph $\mathcal{G}_{\bar{\psi}}$, (lower-left) edge-computing graph $\mathcal{G}_{\Theta}$ and (lower-right) communication graph $\mathcal{G}_{c}$,.
VIII Conclusions